What line is perpendicular to the line y=-1/2x-2?
1
Expert's answer
2020-12-17T19:15:50-0500

"y=mx+b" -- the equation of line.

In our case "m=-\frac{1}{2}, b=2"


Let "y=m_1x+b_1" be the equation of perpendicular line.


Perpendicular line must have a slope that is the negative reciprocal of the original slope.

So "m_1=-\frac{1}{m}=2"


"b_1" can be any real number.


So "y=2x+b_1" is the equation of perpendicular line, where "b_1" can be any real number.